OpenStudy (anonymous):

help need answer to this math proplome! Each year, Terrell buys a plane ticket to visit his grandmother for a week. Last year, he paid one hundred thirty dollars. This year, he paid one hundred forty-three dollars. What was the percent increase in the cost of his plane ticket? Show your work.

OpenStudy (anonymous):

If x is the original cost and y is the new cost, then (y-x)/x times 100 is the percentage increase.

OpenStudy (anonymous):

so just - them and times the left overs by 100 to get the answer

OpenStudy (anonymous):

sort of... percent change is equal to (new price - old price) / (old price) so it would be like (143 - 130)/130

OpenStudy (anonymous):

Don't forget my divisor.

OpenStudy (anonymous):

it's easier to see with small numbers... if you pay $10 today, and $15 tomorrow, the percentage increase is (15-10)/10 = 5/10 = 50%... makes sense, if you think of it... you paid $5 more tomorrow, and $5 is half or 50% of $10.

OpenStudy (anonymous):

ok so 13x 100= 130 so thats the answer

OpenStudy (anonymous):

no, you're missing the step where you divide by the original price.

OpenStudy (anonymous):

What is x and what is y?

OpenStudy (anonymous):

well it is 130 and 143

OpenStudy (anonymous):

Then you can substitute those x and y into the formula I gave you

OpenStudy (anonymous):

ok so - or : ?

OpenStudy (anonymous):

(y-x)/x times 100

OpenStudy (anonymous):

o ok so i get it 143-130 /130 x 100 =

OpenStudy (anonymous):

Yes! You are getting it!

OpenStudy (anonymous):

ok so it = 10%

OpenStudy (anonymous):

Yes! you are done! Good job!

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

for the help

OpenStudy (anonymous):

You are quite welcome! Nice working with you!